OpenStudy (anonymous):

On a cold winter day, a penny (mass 2.50 \rm g) and a nickel (mass 5.00 \rm g) are lying on the smooth (frictionless) surface of a frozen lake. With your finger, you flick the penny toward the nickel with a speed of 2.10 m/s. The coins collide elastically; calculate both their final speeds.

OpenStudy (anonymous):

so i think you just have to use the rule of conserved momentum (and no friction coefficient) and then do some algebra, and plug in values for mass and velocity

OpenStudy (anonymous):

do you do things like velocty initial = velocity final? or something that sounds vaguely familiar?

OpenStudy (anonymous):

i use conservation of kinetic energy and momentum however im not getting the right answer

OpenStudy (anonymous):

can you please post the formula you're using (i can't remember off the top of my head)

OpenStudy (anonymous):

conservation of momentum m1U1 +m2U2 = m1V1 +m2V2 conservation of kinetic energy 1/2)m1U^21 +(1/2)m2U2^2 = (1/2)m1V^21 +(1/2)m2V2^2

OpenStudy (anonymous):

is U velocity?

OpenStudy (anonymous):

initial velocity

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that really meant to be to the power of 21

OpenStudy (anonymous):

the 1 is a subscript and the 2 is the power

OpenStudy (anonymous):

oh ok, so it'd look like \[V _{2}^{2}\] etc

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

ok well i have these two 2.5 * 2.1 + 5 * 0 = 2.5 * V1 + 5 * V2 ------- eqn 1 therefore 2.5*V1 = -5.25+5 * V2 V1 = -2.1 + 2* V2-------eqn2 and V2 = 1.05+0.5 * V1-------eqn 3 can you use that?

OpenStudy (anonymous):

oh no, i hav eto go to work...

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

how did u get equation 3

OpenStudy (anonymous):

rearranging the un numbered one, so add 5.25 to both sides, leaves 5.25+2.5 * V1 = 5* V2 then divide both sides by 5 which leaves 5.25/5 +2.5/5 * V1 = V2

OpenStudy (anonymous):

then you might be able to sub those equations individually into the kinetic formula you get to solve for V1 then do it again to solve for V2... so use eqn 2 in kinetic for V1 and eqn 3 for V2 sorrry i can't help more, i gotta go.

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

have to come back tonight... in about 12 hrs. maybe 11 if im feeling not tired at 11pm

OpenStudy (anonymous):

if you can't get it, i'll try to finish the question with all working on here posted just leave another message or whatever and ill be back.

OpenStudy (anonymous):

What speed did you get initially that was wrong?

OpenStudy (mos1635):

nath aplay conservation of momentum aplay conservation of kinetic energy you now have 2 equations solve system of equetions to V1 and V2... m1*v=M1*V1+m2*V2 M1( v-V1)=M2*V2 (that is eq 1) 1/2m1v^2=1/2m1V1^2+1/2 m2 V2^2 m1 V^2=m1 V1^2+m2 V2^2 m1*(V^2-V1^2)=M2*V2^2 m1*(V-V1)*(V+V1)=m2*V2^2 divide that with (1) and you have..... V+V1=V2 (that is eq 2) solve (1) and (2) to V1 and V2

OpenStudy (anonymous):

or do it that way... whatever. :P